OpenStudy (gabylovesyou):

@sleepyjess what is the equation in point-slope form of the line passing through (2, 5) and (-1, 8)

OpenStudy (uri):

slope formula

OpenStudy (sleepyjess):

We know that point slope form is \(\sf y-y_1 = m(x-x_1)\). Do you know the slope?

OpenStudy (uri):

\[m =\frac{ (y2-y1) }{(x2-x1)? }\]

OpenStudy (gabylovesyou):

-1 ?

OpenStudy (sleepyjess):

\(\sf\dfrac{8-5}{2+1}\)

OpenStudy (uri):

 plug in and solve brah